3. What average force is required to stop a 950 -Kg car in 80 sifthe car is traveling at 95 Km/h?
Answer #1

here,

initial speed , u = 95 km/h

u = 26.39 m/s

time taken , t = 8 s

let the accelration be t

v = u +a * t

0 = 26.39 + a * 8

a= - 3.3 m/s^2

mass , m = 950 kg

the average force required , F = m * a

F = 950 * 3.3 N

F = 3133.7 N